# Flexible game-playing AI with AlphaViT: adapting to multiple games and board sizes

### v1

@@ 6 Conclusion and discussion
We propose AlphaViT, AlphaViD, and AlphaVDA, which are game-playing AI agents designed to address the limitations of AlphaZero using ViT. Unlike AlphaZero, which is limited to fixed board sizes, these proposed methods can effectively handle variations in board size, demonstrating flexibility and adaptability in their gameplay across different board sizes and game types. In addition, we showed that AlphaViT, AlphaViD, and AlphaVDA can simultaneously train and play multiple games, such as Connect4, Gomoku, and Othello, using a single model trained on all games. This ability to generalize across games with a single model represents a significant advance over traditional game-specific AI models.
The results of our experiments show that AlphaViT and AlphaViD outperform baseline methods such as Minimax and MCTS in most configurations. Although AlphaZero still achieves the highest Elo ratings in some cases, especially in games with larger boards, the proposed agents demonstrate competitive performance, especially in Othello, where AlphaViT approaches and even exceeds AlphaZero’s Elo rating in specific configurations. Furthermore, the multi-game versions of AlphaViT and AlphaViD perform on par with their single-game counterparts, further highlighting the ability of these architectures to generalize and adapt across different board sizes.
@@
AlphaViT and AlphaViD showed strong adaptability across different games and board sizes. In particular, agents trained on single games often demonstrate performance comparable to AlphaZero, even when playing on board sizes for which they were not explicitly trained. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes in which skills from simpler game variants (e.g., 9x9 Go) can be applied to more complex versions (19x19 Go). This similarity to human learning patterns suggests that such training paradigms may be beneficial for AI development, particularly in the context of multitask learning.
@@
A comparison of the three proposed architectures shows that AlphaViT performs slightly better than AlphaViD and AlphaVDA despite having the same number of parameters. This may be due to the fewer transformer encoder layers in AlphaViD and AlphaVDA, which rely on a more complex architecture with both encoder and decoder layers. AlphaViT’s simpler architecture, consisting only of encoder layers, may benefit from having more layers dedicated to learning, resulting in more efficient performance in certain games. However, this simplicity limits AlphaViT’s flexibility because its output size is fixed to the number of input tokens, reducing its applicability to games beyond the classic board games tested here. In contrast, AlphaViD’s inclusion of a decoder layer allows it to adjust the size of its policy vector dynamically, providing greater adaptability to games with different action spaces. This architectural flexibility will make AlphaViD more versatile for handling complex games or environments with continuous action spaces.
@@
Future work will explore the application of these architectures to a broader range of games, including those with more complex rules and non-deterministic elements. In addition, we will extend the flexibility of ViT to other deep reinforcement learning methods, such as deep Q-network, and develop a game AI agent that can play more flexibly, including computer games.
@@ 1 AlphaZero
AlphaZero consists of a deep neural network (DNN) and Monte Carlo tree search (MCTS), as shown in Fig. 1. The DNN receives an input representing the current state of the board and the current player. It then outputs the estimated state value and the move probability. MCTS determines the best move based on the value and move probability. AlphaViT, AlphaViD, and AlphaVDA adopt this same fundamental structure, employing an identical decision-making process to select their next moves.
@@ 1.1 Deep neural network in AlphaZero
AlphaZero’s deep neural network (DNN) predicts the value $v(s)$ and the move probability $\bm{p}$ with components $p(a|s)$ for each action $a$, given a state $s$. In a board game context, $s$ and $a$ represent the board state and the move, respectively. The DNN receives an input representing the current board state and the current player’s disc color. Fig. LABEL:fig:network illustrates the DNN architecture, which consists of a Body (residual blocks) and a Head (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}$.
The input to the DNN is an $M\times N\times(2T+1)$ image stack that contains $12T+$ 1 binary feature planes of size $M\times N$. Here, $M\times N$ refers to the board size, and $T$ is the number of histories. The first $T$ feature planes represent the occupancy of the player’s discs, with a feature value of 1 indicating that a disc occupies the corresponding cell, and 0 indicating otherwise. Similarly, the following $T$ feature planes represent the occupancy of the other players’ discs. The last feature plane represents the disc color of the current player, with the disc colors of the first and second players being represented by 1 and -1, respectively.
@@ 1.2 Monte Carlo tree search in AlphaZero
In Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has a maximum score, as described by the following equation:
@@
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time.
In Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{l})$ and $\bm{p}_{a}(s_{l})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,P(s_{L},a)=p_{a}\}$.
In Backup, the visit counts and values are updated in a backward pass through each step, $t\leq L$. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a)\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
@@ 2 Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero use a common training scheme. The training process consists of three main components: Self-play, Augmentation, and Learning, which are iterated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted to allow training on a single computer.
During the Self-play phase, the AI agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected from valid moves based on the softmax policy:
@@
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$ the most visited action is selected. Through Self-play, we collect board states, winners, and search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
In the Augmentation phase, the dataset derived from Self-play is augmented by introducing symmetries specific to the game variant (e.g., two symmetries for Connect4 and eight for Othello and Gomoku). This augmented data are added to a queue with a capacity of $N_{\mathrm{queue}}$ states to form the training dataset.
For the first learning iteration, the training data queue is filled with data generated by self-play using MCTS100 and augmented them. In subsequent iterations, new data generated by Self-play are added to the training data queue. To learn multiple games simultaneously, we prepare a separate training data queue for each game.
During the Learning phase, the deep neural network (DNN) is trained using mini-batch stochastic gradient descent with $N_{\mathrm{batch}}$ batch size and $N_{\mathrm{epochs}}$ epochs. The optimization process includes momentum and weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$:
@@
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Learning phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect4, Gomoku, and Othello, the mini-batch of Connect4, the mini-batch of Gomoku, and the mini-batch of Othello are used orderly for update.
@@ 3 Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are detailed in Table LABEL:tab. The weight decay and momentum values are consistent with those specified in the AlphaZero pseudocode (Silver et al., 2018). All other parameters for AlphaZero are consistent with the previous implementation Fujita (2022). The hyperparameters for the other models were carefully hand-tuned to optimize their performance.
Table LABEL:tab:param_alphavit_game lists the game-specific hyperparameters for AlphaViT, ALphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size. These hyperparameters were also carefully hand-tuned.
@@
Learning rate 0.01 0.01 0.01 0.01
Momentum 0.9 0.9 0.9 0.9
Weight decay 0.0001 0.0001 0.0001 0.0001
@@
stride of a patch 1 1 1 -
Num of encoder layers $L$ $L$ $L$
@@
forward size of encoder 1024 1024 1024 -
Num of encoder head 16 16 16 -
@@
forward size of encoder - 1024 1024 -
Num of decoder head - 16 16 -
Action token size - - 256 -
Num of residual blocks - - - 3
@@ 4 Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game. For the Connect4 variants, the values of the connections of two and three same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$ and $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively, where $R$ is the base reward, and $c_{\mathrm{disc}}$ and $c_{\mathrm{minimax}}$ are the colors of the connecting discs and the minimax player’s disc, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have a value of $R^{3}c_{\mathrm{win}}c_{\mathrm{minimax}}$, where $c_{\mathrm{win}}$ is the color of the winner, and $R=100$.
@@
where $v_{(}x,y)$ is the value of cell $(x,y)$ and $o(x,y)$ is the occupancy of cell $(x,y)$. For 6x6 Othello and 8x8 Othello, the minimax algorithm evaluates each cell using Eq. 12 and 13, respectively. $o(x,y)$ is $1$, $-1$, and 00 if cell $(x,y)$ is occupied by the first player’s disc, the second player’s disc, and empty, respectively. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The terminal nodes have a value of $E_{\mathrm{end}}=1000c_{\mathrm{win}}c_{\mathrm{minimax}}$.

### v2

@@ 6 Conclusion and discussion
In this paper, we present AlphaViT, AlphaViD, and AlphaVDA, which are novel game-playing AI agents designed to overcome the limitations of AlphaZero using Vision Transformers (ViT). Unlike AlphaZero, which is restricted to fixed board sizes, our proposed agents demonstrate adaptability, effectively handle variations in board size, and exhibit flexibility across games. Furthermore, these agents can simultaneously train on and play multiple games, such as Connect4, Gomoku, and Othello, within a single shared neural network. The performance of these multitask agents surpasses that of traditional game AI algorithms and, in some cases, approaches the performance of AlphaZero.
The results of this study show that AlphaViT, AlphaViD, and AlphaVDA outperform traditional methods such as Minimax and MCTS across all tested scenarios. Although AlphaZero remains the top performer in some cases, particularly for games with larger boards, the proposed agents exhibit competitive performance. In Othello, AlphaViT, AlphaViD with a deep DNN (AlphaViD L5), and AlphaVDA with a deep DNN (AlphaVDA L5) achieve Elo ratings equal to or higher than AlphaZero. Furthermore, multigame-trained variants perform on par with or better than single-game-trained variants.
@@
AlphaViT, AlphaViD, and AlphaVDA show strong adaptability across different games and board sizes. The agents with DNNs trained on a single game often achieve performance comparable to traditional game algorithms, such as Minimax and MCTS, even when playing on board sizes on which they are not trained. Moreover, multi-game-trained agents frequently perform on par with or surpass their single-game-trained counterparts. The agents with fine-tuned DNNs trained on small-board games consistently achieve performance equal to or better than that of agents with non-fine-tuned DNNs. In the case of Gomoku, using pre-trained weights from small-board games significantly accelerates convergence and enhances the final performance. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes which skills acquired in simpler game variants (e.g., 9x9 Go) can be applied to more complex versions (e.g., 19x19 Go). Such adaptability suggests that the proposed agents may have significant potential for advancing the development of multitask AI.
@@
Comparing the three proposed agents reveals that AlphaViT L4 outperforms both AlphaViD L1 and AlphaVDA L1, despite having an equivalent number of parameters. This difference in performance may be attributed to the fewer transformer encoder layers in AlphaViD and AlphaVDA. This observation is further supported by the fact that AlphaViD L5 and AlphaVDA L5 exhibit performance comparable to AlphaViT L4. The simpler architecture of AlphaViT, consisting solely of encoder layers, may lead to more efficient performance in certain games, even when it has fewer parameters than its variants. However, this simplicity constrains AlphaViT’s flexibility, as its output size is fixed to the number of input embeddings, limiting its applicability to games beyond the classic board games. In contrast, including a decoder layer in AlphaViD and AlphaVDA allows for dynamic adjustment of the policy vector size, providing greater adaptability to games with varying action spaces. This architectural flexibility makes AlphaViD and AlphaVDA versatile candidates for handling more complex games or environments with continuous action spaces.
Our results also relate to those of Soemers et al. (2023), who highlighted the importance of transferring trained policies and value functions across games with varying board sizes and action spaces. Similarly, our agents, equipped with DNNs that utilize weights fine-tuned from a small-board game or trained simultaneously on multiple games, demonstrate enhanced gameplay skills. This suggests that knowledge obtained from small-board games and other games is used efficiently. The ability to train on multiple games concurrently may also help these agents avoid overfitting, further enhancing their generalization capabilities.
Although our proposed agents with transformer-based DNNs are slightly weaker than AlphaZero, we believe they have the potential for further improvement through enhanced training techniques. For example, an efficient decay of the learning rate during training iterations (Silver et al., 2016, 2016, 2018) and warm-start method (Wang et al., 2020, 2021) will improve performance. Increasing the number of self-plays per iteration may also be effective.
@@
In future work, we plan to extend these architectures to a broader range of games, including those with more complex rules and stochastic elements. In addition, we aim to incorporate the flexibility of ViT-based DNNs into other deep reinforcement learning frameworks, such as deep Q-networks, to create AI agents capable of playing a wider variety of games, including video games, with enhanced adaptability.
@@ 1 AlphaZero
AlphaZero integrates a deep neural network (DNN) with Monte Carlo Tree Search (MCTS), as illustrated in Fig. 1. The DNN processes input representing the current board state and the player, producing an estimated state value and move probabilities. MCTS then uses these outputs to identify the optimal move. Similarly, AlphaViT, AlphaViD, and AlphaVDA adopt this foundational framework.
@@ 1.1 Deep neural network in AlphaZero
AlphaZero’s deep neural network (DNN) predicts the value $v(s)$ and the move probability $\bm{p}$ with components $p(a|s)$ for each action $a$, given a state $s$. In the board game context, $s$ and $a$ represent the board state and the move, respectively. In the context of board games, $s$ represents the board state, and $a$ denotes a move. The DNN receives input representing the current board state and the current player’s disc color. Fig. A1 illustrates the DNN architecture, which consists of a Body (residual blocks) and a Head (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}$.
The input to the DNN is an $H\times W\times(2T+1)$ image stack that contains $12T+$ 1 binary feature planes of size $H\times W$. Here, $H\times W$ refers to the board size, and $T$ is the number of histories. The first $T$ feature planes represent the occupancy of the player’s discs, with a feature value of 1 indicating that a disc occupies the corresponding cell, and 0 otherwise. Similarly, the following $T$ feature planes represent the occupancy of the opponent’s discs. The last feature plane represents the disc color of the current player, with the disc colors of the first and second players represented by 1 and -1, respectively.
Figure A1: Deep neural network of AlphaZero.
@@ 1.2 Monte Carlo tree search in AlphaZero
In the Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has the maximum score, as described by the following equation:
@@
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time. Additionally, when the parent node is the root node, the node selection is performed using an $\epsilon$-greedy algorithm based on the UCT scores.
In the Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{l})$ and $\bm{p}_{a}(s_{l})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,P(s_{L},a)=p_{a}\}$.
In the Backup, the visit counts and values are updated in a backward pass through each step, $t\leq L$. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a)\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in the Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
@@ 2 Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero use the same training scheme. The training process consists of three main components: Self-play, Augmentation, and Update, which are iterated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted to allow training on a single computer.
During the Self-play phase, an agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected from valid moves based on the softmax policy as follows:
@@
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$, the most visited action is selected. Through Self-play, we collect board states, winners, and search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
In the Augmentation phase, the dataset derived from Self-play is augmented by introducing symmetries specific to the game variant (e.g., two symmetries for Connect4 and eight for Othello and Gomoku). This augmented data is added to a queue with a capacity of $N_{\mathrm{queue}}$ states to form the training dataset.
For the first update iteration, the training data queue is filled with data generated by self-play using MCTS100, which is then augmented. In subsequent iterations, new data generated by Self-play are added to the training data queue. To simultaneously learn multiple games, we prepare a separate training data queue for each game.
During the Update phase, the deep neural network (DNN) is trained using mini-batch stochastic gradient descent with $N_{\mathrm{batch}}$ batch size and $N_{\mathrm{epochs}}$ epochs. The optimization process includes momentum and weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$:
@@
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Update phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect4, Gomoku, and Othello, a mini-batch of Connect4, a mini-batch of Gomoku, and a mini-batch of Othello are used orderly for the update.
@@ 3 Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are detailed in Table A1. The weight decay and momentum values are consistent with those specified in the AlphaZero pseudocode (Silver et al., 2018). All other parameters for AlphaZero are consistent with the previous implementation (Fujita, 2022). The hyperparameters of the other models were carefully hand-tuned to optimize their performance.
Table A2 lists the game-specific hyperparameters for AlphaViT, ALphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size. For AlphaViT, AlphaViD, and AlphaVDA, we employ the AdamW optimizer in PyTorch, with all parameters set to their default values except for the learning rate. These hyperparameters were also carefully hand-tuned.
@@
Learning rate 0.0001 0.0001 0.0001 0.05
Momentum - - - 0.9
Weight decay 0.01 0.01 0.01 0.0001
@@
Stride for making patches 1 1 1 -
@@
Dim feedforward of encoder 1024 1024 1024 -
Num of encoder heads 8 8 8 -
Size of positional embeddings $512\times 256$ $512\times 256$ $512\times 256$
@@
Dim feedforward of decoder - 1024 1024 -
Num of decoder head - 8 8 -
Size of action embeddings - - 256 -
Dropout rate 0.1 0.1 0.1 -
Num of residual blocks - - - 6
@@
4 Monte Carlo tree search
MCTS is a best-first search method that does not require a predefined evaluation function (Winands, 2017). It operates through four strategic steps: Selection step, Playout step, Expansion step, and Backpropagation step (Winands, 2017). In the Selection step, the tree is traversed from the root node to a leaf node. Child node $c$ of parent node $p$ is selected to maximize the following score:
$\mathrm{UCT}=q_{c}/N_{c}+C\sqrt{\frac{\ln(N_{p}+1)}{N_{c}+\varepsilon}},$ (12)
where $q_{c}$ is the cumulative value of $c$, $N_{p}$ is the visit count of $p$, $N_{c}$ is the visit count of $c$, and $\varepsilon$ is a constant value to avoid division by zero. In the Playout step, a valid move is selected at random until the end of the game is reached. In the Expansion step, child nodes are added to the leaf node when the visit count of the leaf node reaches $N_{\mathrm{open}}$. In the Backpropagation step, the result of the playout is propagated along the path from the leaf node to the root node. If the MCTS player itself wins, loses, and draws, $q_{i}\leftarrow q_{i}+1$, $q_{i}\leftarrow q_{i}-1$, and $q_{i}\leftarrow q_{i}$, respectively. $q_{i}$ is the cumulative value of node $i$. The sequence of these four steps constitutes a single simulation. The simulation process is repeated $N_{\mathrm{sim}}$ times. After all simulations, MCTS selects the action corresponding to the most visited child node of the root. In this study, $C=0.5$, $\varepsilon=10^{-7}$, and $N_{\mathrm{open}}=5$.
@@ 5 Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game.
For the Connect4 variants, the values of the connections of two and three same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$ and $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively, where $R$ is the base reward, and $c_{\mathrm{disc}}$ and $c_{\mathrm{minimax}}$ are the colors of the connecting discs and the minimax player’s disc, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have a value of $R^{3}c_{\mathrm{win}}c_{\mathrm{minimax}}$, where $c_{\mathrm{win}}$ is the color of the winner, and $R=100$.
@@
where $v_{(}x,y)$ is the value of cell $(x,y)$ and $o(x,y)$ is the occupancy of cell $(x,y)$. For 6x6 Othello and 8x8 Othello, the minimax algorithm evaluates each cell using Eq. 14 and 15, respectively. $o(x,y)$ is $1$, $-1$, and 00 if cell $(x,y)$ is occupied by the first player’s disc, the second player’s disc, and empty, respectively. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The terminal nodes have a value of $E_{\mathrm{end}}=1000c_{\mathrm{win}}c_{\mathrm{minimax}}$.